OpenStudy (s):

c(x) = 6x^2 + 14x + 18 a) find marginal cost function b) the cost and marginal cost when 16 units are produced c) cost of producing the 16th unit

OpenStudy (anonymous):

take derivative with respect to x and set it equal to zero then solve for x

OpenStudy (s):

12x + 14 = 0 12x = -14 x = -14/12 ?

OpenStudy (s):

What is the marginal cost function? What is marginal cost?

OpenStudy (anonymous):

The marginal cost function is the derivative of the marginal cost function. So take the derivative of c(x) \[\huge d\frac{c(x)}{dx}\]

OpenStudy (anonymous):

c'(x)=12x+14

OpenStudy (s):

Oo ok thank you. So that is it. How about part b? Do i just plug in 16 for x? c'(x)=12x+14= 12(16) + 14 = 206 ??? And how would i answer part c ?

OpenStudy (anonymous):

In part b for the cost substitute 16 for x in the cost function then substitute 16 for x in the marginal cost function or the derivative of c(x) which you have done. for just add the marginal coast to the cost. so c(16)+c'(x) evaluated @ 16 is the cost of producing 16units. \[\huge c(16)+c'(16)\]

OpenStudy (anonymous):

I meant to write for c just add the marginal coast to the cost. so c(16)+c'(x) evaluated @ 16 is the cost of producing 16units.

OpenStudy (s):

Ooo ok I got thank you very very much!

OpenStudy (s):

got it*